# The orbital evolution of the tidally stripped star and disk-driven stable mass transfer for QPEs in GSN 069

### v4

@@ 2.5 Comparison with the observation of QPEs
Before QPEs from GSN 069 disappeared, its peak intensity in 0.4–1 keV decays dramatically (Miniutti et al., 2023b). However, the bolometric luminosity by assuming absorbed black-body spectrum was almost constant for each observation, except for the last one, where it was reduced by roughly half (Miniutti et al., 2023a). If the bolometric luminosity is positively correlated with the mass loss of the star at the periastron, such result implies that $\beta$ is constant and then gets smaller.
Assuming the QPEs is caused by the mass transfer via Roche-lobe overflow, the orbiting star should be a WD for GSN 069 (King, 2020). As shown in Fig. 1, the mass loss for WD grows rapidly except in the final stage of evolution, when the orbital period also grows rapidly. There was no significant change in the period of QPEs in GSN 069. So a mechanism for stabilizing mass transfer is required to explain the long-term evolution of QPEs, and the drag of a gaseous disk may achieve this.
@@ 3.1 Drag force by the accretion disk
The drag force caused by the accretion disk can be expressed as
@@
Here $\mathbf{v}_{rel}=\mathbf{v}-\mathbf{v}_{k}$ is the relative velocity of the star with respect to the local gas, where $\mathbf{v}$ is the speed of star and the keplerian velocity $\mathbf{v}_{k}$ is the velocity of the local gas. The timescale $\tau_{F}$ of the drag force, which is considered as hydrodynamic force or dynamical friction. The hydrodynamic timescale is (Szölgyén et al., 2022)
@@
Here $C_{D}$ is related to the shape of the star, and equals 1 for a sphere, and $\rho_{g}$ is the density of the local gas. The timescale of the dynamical friction is (Ostriker, 1999)
@@
Here $\mathcal{M}=v_{\mathrm{rel}}/c_{s}$ is the Mach number where $c_{s}$ is the local sound speed. And $r$ is the characteristic size, and we adopted the distance from the central black hole.
The change of the specific angular momentum $h=\sqrt{G(M_{h}+M)a(1-e)}$ caused by the drag force can be obtained by
@@
where $f$ is the true anomaly, and $\chi$ is 1 for the prograde star and -1 for the retrograde star. And the change of the specific energy $K=-G(M_{h}+M)/2a$ is
@@
The secular evolution can be obtained by averaging them over an orbital period. Then the secular evolution of orbital parameters are
@@
While near the periastron, the velocity of the orbiting star is greater than the Keplerian velocity and is therefore slowed down, as it moves away from the periastron, its velocity becomes less than the Keplerian velocity and is therefore accelerated. Since the latter interaction time is longer, star will likely gain angular momentum from the accretion disk and thus reduce the $\beta$.
Eq. (23) can also be used to calculate instantaneous $\dot{\beta}/\beta$ for different $f$, as shown in Fig. 4. Even for a star near the periastron, $\dot{\beta}$ is still positive, so the drag force by the disk can always suppress mass transfer.
Figure 4: $\dot{\beta}/\beta$ with true anomaly for different eccentricity. For all eccentricity, $\dot{\beta}/\beta$ is always negative and equal to zero while $f=0$.
3.2 Application on GSN 069
@@ 3.2.1 Stable mass transfer
Figure 5: Timescales with true anomaly for the hydrodynamic force and dynamical friction for a WD with $M=0.21M_{\odot}$, $P_{b}=9\mathrm{\leavevmode\nobreak\ hr}$ and $e=0.95$. The upper panel is for the prograde star and the lower panel is for the retrograde star. Solid lines and dashed lines represent the timescales of the hydrodynamic force and dynamical friction, respectively. Different colors represent the different accretion rate of the disk.
@@
We consider an $\alpha$ standard disk with $\alpha=0.1$ and $M_{h}=5\times 10^{5}\mathrm{\leavevmode\nobreak\ M_{\odot}}$. The density and sound speed can be calculated for a given accretion rate $\dot{M}_{Acc}$ in the unit of Eddington accretion rate $\dot{M}_{Edd}$ (radiation efficiency $\eta=0.1$). As shown in Fig. 5, the hydrodynamic force is always stronger than the dynamical friction in all orbital phases, thus dominant on the effect by disk. We only consider the hydrodynamic force in the following calculation.
@@
Figure 6: Secular change of orbital parameters caused by hydrodynamic force for a prograde WD with $M=0.21M_{\odot}$ and $P_{b}=9\mathrm{\leavevmode\nobreak\ hr}$ under different accretion rate. The top panel shows the change of $\beta$, specific angular momentum, and specific orbital energy caused by drag force, it also shows the effect on $\beta$ caused by gravitational radiation for comparison. The dashed black line represents the point with $\dot{M}_{Acc}$ of GSN 069 and $e=0.95$, it predicts $\dot{\beta}/\beta\approx-3.5\times 10^{-4}\mathrm{\leavevmode\nobreak\ yr}^{-1}$(see texts). The bottom panel represents the change of the $a$ and $e$ relative to $\beta$.
@@
Figure 7: As in Fig. 6, but for a retrograde WD. The sign of $\dot{\beta}/\beta$, $\dot{h}/h$, $(\dot{a}/a)/(\dot{\beta}/\beta)$, and $\dot{e}/(\dot{\beta}/\beta)$ is changed. To show value of $-\dot{e}/(\dot{\beta}/\beta)$ in the logarithmic axis, we add 1 on them, because they change sign at high $\dot{M}_{Acc}$.
@@
The secular evolution of orbital parameters is shown in Fig. 6. For varies $\dot{M}_{Acc}$ and $e$, a prograde WD always gains angular momentum and loses energy from the accretion disk, which decreases $\beta$, $a$ and $e$. A retrograde WD shows different evolution, as shown in Fig. 7. The angular momentum is transferred from it to the accretion disk, which makes $\beta$ increase. The energy is still transferred from the WD to the accretion disk. They make $\beta$ increase and $a$ still decrease. And $e$ continues to decrease at low $\dot{M}_{Acc}$ but increases at high $\dot{M}_{Acc}$.
@@
The different dependence of the parameter evolution on $\dot{M}_{Acc}$ at high and low $\dot{M}_{Acc}$ is because the inner region of the accretion disk is dominated by radiation pressure at high $\dot{M}_{Acc}$ and gas pressure at low $\dot{M}_{Acc}$. Therefore, the disk density decreases with increasing $\dot{M}_{Acc}$ for radiation pressure dominance resulting in a decrease in drag force, while the disk density increases with increasing $\dot{M}_{Acc}$ for gas pressure dominance increasing drag force.
@@
A prograde WD is preferred for GSN 069. The WD may be captured from a binary system by the Hill mechanism (Hills, 1988; Wang et al., 2022). The accretion disk is likely to be a fallback disk from a TDE, which is possibly a disruption of the common envelope of the binary system (Wang, 2023). In this scenario, the WD is prograde.
@@
Here we adopt the results in Miniutti et al. (2023b). Assuming radiation efficiency $\eta=0.1$, the accreted mass is about $3.7\times 10^{-8}\mathrm{\leavevmode\nobreak\ M_{\odot}}$, corresponding to $\dot{M}/M\approx-1.7\times 10^{-4}\mathrm{yr}^{-1}$ and $\beta\approx 0.5005$ with $M=0.21\mathrm{\leavevmode\nobreak\ M_{\odot}}$. The change of $\beta$ induced by mass transfer is $\langle\dot{\beta}/\beta\rangle_{\dot{M}}\approx 1.1\times 10^{-4}\mathrm{\leavevmode\nobreak\ yr}^{-1}$ with $\zeta=-1/3$. The bolometric luminosity of TDE is about $10^{44}\mathrm{\leavevmode\nobreak\ erg/s}$ corresponding to $\sim 1.6\dot{M}_{Edd}$. Assuming $e=0.95$, it predicts $\langle\dot{\beta}/\beta\rangle_{disk}\approx-3.5\times 10^{-4}\mathrm{\leavevmode\nobreak\ yr}^{-1}$ as shown in Fig. 6. The effects of disk drag and mass transfer on $\beta$ are in the same order of magnitude, suggesting that disk drag has the ability to stabilize mass transfer.
@@ 3.2.2 History and future of evolution
Early in the TDE, there is a high accretion rate of the fallback disk, resulting in weak drag force by the disk. The evolution of $\beta$ is dominant by the mass transfer or gravitational wave radiation. Thus, the intensity of QPEs kept increasing during this stage. QPEs can be observed when the intensity of QPEs is higher than TDE. It is consistent with the fact that QPEs in GSN 069 do not accompany the TDE at the same time, but appear suddenly during the decay of the TDE (Miniutti et al., 2023b).
As the accretion rate decreases, drag force by the disk begins to dominate $\dot{\beta}$, and at that point the intensity of the QPEs begins to decrease. When the TDE rebrighten, the increased accretion rate led to a weaker drag force and thus an increase in the mass transfer. Then, when the accretion rate begins to decrease again, the effect of the drag force enhances and the intensity of QPEs will keep decreasing. After switching from radiation pressure dominance to gas pressure dominance in the inner region of the disk, the drag force decreases with decreasing accretion rate. When $\dot{\beta}$ is dominated by gravitational wave radiation, the mass transfer begins to enhance. When $\dot{\beta}$ is dominated by mass transfer, the unstable mass transfer leads to WD escape from the SMBH.
While the accretion rate decreases, the outer radius of the fallback disk may change (Shen & Matzner, 2014). The orbital semi-long axis of the stripped WD may be larger than the outer radius of the disk. In this situation, the drag force by the disk acts only on that side of the periastron. The secular evolution of the orbital parameters can be obtained by considering drag force only at $-\kappa\pi<f<\kappa\pi$ when averaging them over an orbit. Here $\kappa$ ranges from 0 to 1. As shown in Fig. 8, the secular evolution of orbital parameters is significantly affected by $\kappa$, especially for $\dot{\beta}/\beta$. When $\kappa<0.5$, the drag force only acts near the periastron, thus the angular momentum of the stripped WD is transferred to the disk as expected.
Figure 8: Distribution of secular change of orbital parameters for different $\dot{M}_{Acc}$ and $\kappa$. The top, middle and bottom panels represent $\dot{h}/h$,$\dot{K}/K$ and $\dot{\beta}/\beta$, respectively. They are obtained with the drag force acting only when true anomaly $f$ is within $-\kappa\pi$ to $\kappa\pi$. The solid line in the top panel represents the $\dot{h}=0$.
Even though the change in the outer radius of the disk affects the evolution of $\beta$, the tendency for the effect of the drag force by the disk to diminish is constant. Therefore, $\beta$ will start to grow when the effect of gravitational wave radiation exceeds the drag force by the disk. Finally, the stripped WD will escape from SMBH after entering self-unstable mass transfer, as described by Section 2.
@@ 3.2.3 Secular evolution of GSN 069
Due to the evolution on the accretion rate of the fallback disk in GSN 069, the secular evolution on mass transfer may experience complex changes. Fig. 9 show an example with possible parameters of GSN 069. Following the formula in Miniutti et al. (2023b), the accretion rate of each peak for TDE is described by
@@
where the parameters can be obtained by fitting the light curve (Miniutti et al., 2023b). For the first peak, $t_{0}=7.6$ years and $t_{\text{peak }}=600$ days, only the decay phase is calculated. For the second peak, $t_{0}=4.8$ year, $t_{\text{peak }}=3850$ days and $\sigma=400$ days. And $n$ is -9/4 for each peak, $\dot{M}_{\text{peak }}$ is 9 $\dot{M}_{Edd}$ for the first peak and 6.3 $\dot{M}_{Edd}$ for the second peak. The orbital parameters are same as them in Section 3.2.1, while the time is setting to $7$ years. Then the past and future evolution can be obtained by the same method in Section 2.3.
@@
Figure 9: Possible evolution of mass transfer with the evolution of the accretion rate of the accretion disk. Solid line and dashed line represent the evolution of the mass transfer at the periastron and the accretion rate of the disk in GSN 069, respectively. See text for details of calculations.
@@
It should be stressed that the results in Fig. 9 is very sensitive to the magnitude of parameters, mainly $\beta$ and $\dot{M}_{\text{peak }}$. The former affects the magnitude of $\dot{M}$ as it enlarges by a factor of $1/(\beta-\beta_{0})$ compared to the $\dot{\beta}$, as shown in Eq. 3. The latter, on the other hand, affects the magnitude of the disk drag, thus changing the moment of transition when $\beta$ changes from increasing to decreasing.
@@
We do not show the evolution of the light curve directly, because the radiative processes are not considered here. If one assumes a positive correlation between the bolometric luminosity and the accretion mass, the disappearance of the QPEs in the early stages of the TDE and the lack of increase of its luminosity as well as its fading can be explained. That is, the early accreted mass increases rapidly and then starts to decrease as the disk drag increases.
@@
However, the subsequent evolution is somewhat difficult to explain. Although the decrease in the accreted mass when the TDE flares again seems to explain the disappearance of the QPEs. But as pointed out by Miniutti et al. (2023a), the spectrum of QPEs need to be colder than it of the TDE to be undetectable, not just less luminous than the TDE. In addition, the second flare of the two QPEs that reappeared brightened, as expected in Fig. 9, yet the first flare did not brighten and could not be explained by our model.
@@
These variations may be related to radiation processes of the QPEs. Lu & Quataert (2023) suggested that the QPEs is produced by the circularized shock of the accreted matter moving inside the disk. The peak temperature decreases during the second outburst of the TDE due to the decrease in disk density. This may explain the disappearance of QPEs. Thus the evolution of QPEs is also directly related to the evolution of accretion disk. The secular evolution on the light curve and spectrum of QPEs requires a further analysis combined with radiative processes in the future.
@@
For the accretion model of the QPEs, the long-short behavior on the intensity and recurrence time is difficult to explain. King (2022) proposed that the short accretion timescale may cause the WD to fail to maintain hydrostatic equilibrium and oscillate in radius, which may explain this feature of QPEs. Such oscillations may be perturbed when the density of the surrounding gas changes, which may explain the disappearance of such features in QPEs before the second outburst of the TDE and the weaker QPE after this outburst that did not enhance as we expected. The evolution of such short timescales is beyond the scope of this paper, which focuses on long-term evolution.

### v5

@@ 2.5 Comparison with the observation of QPEs
Before the QPEs from GSN 069 disappeared, its peak intensity in 0.4–1 keV decayed dramatically (Miniutti et al., 2023b). However, the bolometric luminosity by assuming absorbed black-body spectrum was almost constant for each observation, except for the last one, where it was reduced by roughly half (Miniutti et al., 2023a). If the bolometric luminosity is positively correlated with the mass loss of the star at the periastron, such result implies that $\beta$ is constant and then gets smaller.
Assuming the QPEs is caused by the mass transfer via Roche-lobe overflow, the orbiting star should be a WD for GSN 069 (King, 2020). As shown in Fig. 1, the mass loss for WD grows rapidly except in the final stage of evolution, when the orbital period also grows rapidly. There was no significant change in the period of QPEs in GSN 069. Thus, a mechanism for stabilizing mass transfer would required to explain the long-term evolution of QPEs and the drag of a gaseous disk may achieve this effect.
@@ 3.1 Drag force by the accretion disk
The drag force caused by the accretion disk can be expressed as:
@@
Here, $\mathbf{v}_{rel}=\mathbf{v}-\mathbf{v}_{k}$ is the relative velocity of the star with respect to the local gas, where $\mathbf{v}$ is the speed of star and the keplerian velocity, $\mathbf{v}_{k}$, is the velocity of the local gas. The timescale, $\tau_{F}$, of the drag force, which is considered as hydrodynamic force or dynamical friction. The hydrodynamic timescale is (Szölgyén et al., 2022):
@@
Here, $C_{D}$ is related to the shape of the star and is equal to 1 for a sphere, while $\rho_{g}$ is the density of the local gas. The timescale of the dynamical friction is (Ostriker, 1999):
@@
Here, $\mathcal{M}=v_{\mathrm{rel}}/c_{s}$ is the Mach number, where $c_{s}$ is the local sound speed. Also, $r$ is the characteristic size and we adopted the distance from the central black hole.
The change in the specific angular momentum, $h=\sqrt{G(M_{h}+M)a(1-e)}$, caused by the drag force can be obtained via:
@@
where $f$ is the true anomaly, while $\chi$ is 1 for the prograde star and -1 for the retrograde star. And the change in the specific energy $K=-G(M_{h}+M)/2a$ is:
@@
The secular evolution can be obtained by averaging them over an orbital period. Then the secular evolution of orbital parameters are:
@@
When near to the periastron, the velocity of the orbiting star is greater than the Keplerian velocity and it is thus slowed down as it moves away from the periastron. When its velocity becomes lower than the Keplerian velocity, it is therefore accelerated. Since the latter interaction time is longer, the star will likely gain angular momentum from the accretion disk and thus reduce $\beta$.
Equation (23) can also be used to calculate instantaneous $\dot{\beta}/\beta$ for different values of $f$, as shown in Fig. 4. Even for a star near the periastron, $\dot{\beta}$ is still positive, so the drag force by the disk may always suppress the mass transfer.
Figure 4: $\dot{\beta}/\beta$ with the true anomaly for a different eccentricity. For all eccentricities, $\dot{\beta}/\beta$ is always negative and equal to zero, while $f=0$.
3.2 Application to GSN 069
@@ 3.2.1 Stable mass transfer
Figure 5: Timescales with the true anomaly for the hydrodynamic force and dynamical friction for a WD with $M=0.21M_{\odot}$, $P_{b}=9\mathrm{\leavevmode\nobreak\ hr,}$ and $e=0.95$. Upper panel refers to the prograde star and the lower panel to the retrograde star. Solid lines and dashed lines represent the timescales of the hydrodynamic force and dynamical friction, respectively. Different colors represent the different accretion rate of the disk.
@@
We consider an $\alpha$ standard disk with $\alpha=0.1$ and $M_{h}=5\times 10^{5}\mathrm{\leavevmode\nobreak\ M_{\odot}}$. The density and sound speed can be calculated for a given accretion rate, $\dot{M}_{Acc}$, in the unit of Eddington accretion rate, $\dot{M}_{Edd}$ (radiation efficiency $\eta=0.1$). As shown in Fig. 5, the hydrodynamic force is always stronger than the dynamical friction in all orbital phases, thus it dominates the effects from the disk. We only consider the hydrodynamic force in the following calculation.
@@
Figure 6: Secular change in the orbital parameters caused by hydrodynamic force for a prograde WD with $M=0.21M_{\odot}$ and $P_{b}=9\mathrm{\leavevmode\nobreak\ hr}$ under different accretion rate. Top panel shows the change in $\beta$, the specific angular momentum, and the specific orbital energy caused by drag force. It also shows the effect on $\beta$ caused by gravitational radiation for comparison. The dashed black line represents the point with $\dot{M}_{Acc}$ of GSN 069 and $e=0.95$, it predicts $\dot{\beta}/\beta\approx-3.5\times 10^{-4}\mathrm{\leavevmode\nobreak\ yr}^{-1}$(see texts). Bottom panel represents the change in $a$ and $e$ relative to $\beta$.
@@
Figure 7: As in Fig. 6, but for a retrograde WD. The signs of $\dot{\beta}/\beta$, $\dot{h}/h$, $(\dot{a}/a)/(\dot{\beta}/\beta)$, and $\dot{e}/(\dot{\beta}/\beta)$ are changed. To show a value of $-\dot{e}/(\dot{\beta}/\beta)$ in the logarithmic axis, we add 1 to them because they change sign at a high $\dot{M}_{Acc}$.
@@
The secular evolution of orbital parameters is shown in Fig. 6. For varying $\dot{M}_{Acc}$ and $e$, a prograde WD always gains angular momentum and loses energy from the accretion disk, which decreases $\beta$, $a,$ and $e$. A retrograde WD shows different evolution, as shown in Fig. 7. The angular momentum is transferred from it to the accretion disk, which makes $\beta$ increase. The energy is still transferred from the WD to the accretion disk. All this makes $\beta$ increase and $a$ continue to decrease, while $e$ continues to decrease at low $\dot{M}_{Acc}$ but increases at high $\dot{M}_{Acc}$.
@@
The different dependence of the parameter evolution on $\dot{M}_{Acc}$ at high and low $\dot{M}_{Acc}$ is attributed to the inner region of the accretion disk being dominated by radiation pressure at high $\dot{M}_{Acc}$ and gas pressure at low $\dot{M}_{Acc}$. Therefore, the disk density decreases with increasing $\dot{M}_{Acc}$ for radiation-pressure dominance resulting in a decrease in drag force, while the disk density increases with increasing $\dot{M}_{Acc}$ for the gas-pressure dominance increasing the drag force.
@@
A prograde WD is preferred for GSN 069. The WD may be captured from a binary system by the Hill mechanism (Hills, 1988; Wang et al., 2022). The accretion disk is likely to be a fallback disk from a TDE, which is possibly a disruption of the common envelope of the binary system (Wang, 2024). In this scenario, the WD is prograde.
@@
Here, we adopt the results in Miniutti et al. (2023b). Assuming a radiation efficiency of $\eta=0.1$, the accreted mass is about $3.7\times 10^{-8}\mathrm{\leavevmode\nobreak\ M_{\odot}}$, corresponding to $\dot{M}/M\approx-1.7\times 10^{-4}\mathrm{yr}^{-1}$ and $\beta\approx 0.5005$ with $M=0.21\mathrm{\leavevmode\nobreak\ M_{\odot}}$. The change in $\beta$ induced by mass transfer is $\langle\dot{\beta}/\beta\rangle_{\dot{M}}\approx 1.1\times 10^{-4}\mathrm{\leavevmode\nobreak\ yr}^{-1}$ with $\zeta=-1/3$. The bolometric luminosity of the TDE is about $10^{44}\mathrm{\leavevmode\nobreak\ erg/s}$ corresponding to $\sim 1.6\dot{M}_{Edd}$. Assuming $e=0.95$, it predicts $\langle\dot{\beta}/\beta\rangle_{disk}\approx-3.5\times 10^{-4}\mathrm{\leavevmode\nobreak\ yr}^{-1}$, as shown in Fig. 6. The effects of disk drag and mass transfer on $\beta$ are on the same order of magnitude, suggesting that disk drag has the ability to stabilize mass transfer.
@@ 3.2.2 History and future of evolution
Early on in the TDE, there is a high accretion rate of the fallback disk, resulting in weak drag force by the disk. The evolution of $\beta$ is dominated by the mass transfer or gravitational wave radiation. Thus, the intensity of QPEs kept increasing during this stage. Overall, QPEs can be observed when the intensity of QPEs is higher than that of the TDE. This is consistent with the fact that QPEs in GSN 069 do not accompany the TDE at the same time, but appear suddenly during the decay of the TDE (Miniutti et al., 2023b).
As the accretion rate decreases, drag force by the disk begins to dominate $\dot{\beta}$ and, at that point, the intensity of the QPEs begins to decrease. When the TDE rebrightens, the increased accretion rate led to a weaker drag force and, thus, an increase in the mass transfer. Then, when the accretion rate begins to decrease again, the effect of the drag force enhances and the intensity of QPEs will keep decreasing. After switching from radiation-pressure dominance to gas-pressure dominance in the inner region of the disk, the drag force decreases with a decreasing accretion rate. When $\dot{\beta}$ is dominated by gravitational wave radiation, the mass transfer becomes enhanced. When $\dot{\beta}$ is dominated by mass transfer, the unstable mass transfer leads to the WD escaping from the SMBH.
While the accretion rate decreases, the outer radius of the fallback disk may change (Shen & Matzner, 2014). The orbital semi-long axis of the stripped WD may be larger than the outer radius of the disk. In this situation, the drag force by the disk acts only on that side of the periastron. The secular evolution of the orbital parameters can be obtained by considering drag force only at $-\kappa\pi<f<\kappa\pi$ when averaging them over an orbit. Here, $\kappa$ ranges from 0 to 1. As shown in Fig. 8, the secular evolution of orbital parameters is significantly affected by $\kappa$, especially for $\dot{\beta}/\beta$. When $\kappa<0.5$, the drag force only acts near the periastron; thus, the angular momentum of the stripped WD is transferred to the disk, as expected.
Figure 8: Distribution of secular change in orbital parameters for different $\dot{M}_{Acc}$ and $\kappa$. Top, middle, and bottom panels represent $\dot{h}/h$,$\dot{K}/K,$ and $\dot{\beta}/\beta$, respectively. They are obtained with the drag force acting only when true anomaly $f$ is within $-\kappa\pi$ to $\kappa\pi$. The solid line in the top panel represents $\dot{h}=0$.
Even though the change in the outer radius of the disk affects the evolution of $\beta$, the tendency for the effect of the drag force by the disk to diminish is constant. Therefore, $\beta$ will start to grow when the effect of gravitational wave radiation exceeds the drag force by the disk. Finally, the stripped WD will escape from SMBH after entering a self-unstable mass transfer, as described in Section 2.
@@ 3.2.3 Secular evolution of GSN 069
Due to the evolution on the accretion rate of the fallback disk in GSN 069, the secular evolution on mass transfer may experience complex changes. Figure 9 shows an example with possible parameters of GSN 069. Following the formula in Miniutti et al. (2023b), the accretion rate of each peak for the TDE is described by:
@@
where the parameters can be obtained by fitting the light curve (Miniutti et al., 2023b). For the first peak, $t_{0}=7.6$ years and $t_{\text{peak }}=600$ days, only the decay phase is calculated. For the second peak, $t_{0}=4.8$ year, $t_{\text{peak }}=3850$ days and $\sigma=400$ days. And $n$ is -9/4 for each peak, $\dot{M}_{\text{peak }}$ is 9 $\dot{M}_{Edd}$ for the first peak and 6.3 $\dot{M}_{Edd}$ for the second peak. The orbital parameters are same as those given in Section 3.2.1, while the time is set to $7$ years. As a result, the past and future evolution can be obtained using the same method presented in Section 2.3.
@@
Figure 9: Possible evolution of mass transfer with the evolution of the accretion rate of the accretion disk. Solid line and dashed lines represent the evolution of the mass transfer at the periastron and the accretion rate of the disk in GSN 069, respectively. See text for details of the calculations.
@@
It should be stressed that the results in Fig. 9 are very sensitive to the magnitude of parameters, mainly $\beta$ and $\dot{M}_{\text{peak }}$. The former affects the magnitude of $\dot{M}$ as it enlarges by a factor of $1/(\beta-\beta_{0})$ compared to the $\dot{\beta}$, as shown in Eq. 3. The latter, on the other hand, affects the magnitude of the disk drag, thus changing the moment of transition when $\beta$ changes from increasing to decreasing.
@@
We do not show the evolution of the light curve directly, because the radiative processes are not considered here. If we assume a positive correlation between the bolometric luminosity and the accretion mass, the disappearance of the QPEs in the early stages of the TDE and the lack of increase of its luminosity (as well as its fading) can be explained. That is to say, the early accreted mass increases rapidly and then starts to decrease as the disk drag increases.
@@
However, the subsequent evolution is somewhat difficult to explain. Although the decrease in the accreted mass when the TDE flares again seems to explain the disappearance of the QPEs, as pointed out by Miniutti et al. (2023a), the spectrum of QPEs need to be colder than it of the TDE to be undetectable – and not just less luminous than the TDE. In addition, the second flare of the two QPEs that reappeared brightened (as expected from Fig. 9), however, the first flare did not brighten and could not be explained by our model.
@@
These variations may be related to the radiation processes within the QPEs. Lu & Quataert (2023) suggested that the QPEs is produced by the circularized shock of the accreted matter moving inside the disk. The peak temperature decreases during the second outburst of the TDE due to the decrease in disk density. This may explain the disappearance of QPEs. Thus, the evolution of QPEs is also directly related to the evolution of accretion disk. The secular evolution on the light curve and spectrum of QPEs requires a further analysis combined with radiative processes in the future.
@@
For the accretion model of the QPEs, the long-short behavior on the intensity and recurrence time is difficult to explain. King (2022) proposed that the short accretion timescale may cause the WD to fail to maintain hydrostatic equilibrium and oscillate in radius, which may explain this feature of QPEs. Such oscillations may be perturbed when the density of the surrounding gas changes, which may explain the disappearance of such features in QPEs before the second outburst of the TDE and the weaker QPE after this outburst that did not become enhanced, as we expected. The evolution of such short timescales is beyond the scope of this paper, which focuses on long-term evolution.